OpenStudy (jmartinez638):

Find the points of intersection of the graph of the following equations: y = (sqrt3/3x) + 1 x^2 + (y-1)^2 = 4

OpenStudy (anonymous):

\[y=\frac{\sqrt 3}{3x}+1\]?

OpenStudy (jmartinez638):

That is, \[y = \frac{ \sqrt3 }{ 3 }x + 1 \] and \[x^2 + (y-1)^2 = 4\]

OpenStudy (mathmate):

Hints: Given: \(y = \frac{ \sqrt3 }{ 3 }x + 1\) substitute into the second equation: \(x^2 + (y-1)^2 = 4\) \(x^2 + ( \frac{ \sqrt3 }{ 3 }x + 1-1)^2 = 4\) Simplify and solve for x (two roots, symm. about y-axis)